Team:Caltech/Week 9

From 2010.igem.org


iGEM 2010



Home

People

Project Details

Protocols

Completed Systems

Notebook

Biosafety

Human Impact

References

Support


Caltech logo watermark.png


Monday 8/9

  • Group meeting!
  • Re-ran two gels of CPCR results
    • Successful: 55-5, 56-1/-2/-3, 57-1/-2/-3, 54-5/-6
    • Prepped DNA for use in digestions tomorrow

Tuesday 8/10

  • Digested DNA for ligations:
    • L31: L17 + BBa_K215000 + Kan
    • L32: L19 + L20 + Tet
    • L33: L19 + L18 + Tet
    • L34: L17 + L21 + Amp
    • L35: L18 + L21 + Amp
    • L41: L26 + BBa_B0015 + Tet
    • L49: HSP + L40 + Tet
    • L60: BBa_K215000 + L57 + Tet
    • L61: BBa_K215000 + L56 + Tet
    • L63: L29 + L20 + Amp
    • L64: L27 + L20 + Amp
  • Prepped 37-1, 54-5/-6, 55-5, 56, 57 DNA
  • Started CPCR of L36, L38
    • All colonies failed
  • Started liquid culture of BBa_I15008

Wednesday 8/11

  • BBa_I15008 culture did not grow; will re-pick more tomorrow.
  • L43 plate had no colonies; will redo L43 tomorrow.
  • Began CPCR of all successful ligations (31, 32, 33, 34, 35, 36, 38, 41, 49, 60, 61, 62, 63, 64).
  • Gels:
    • gel 1: 35-2 & 63-3 successful
    • gel 2: 64-1 successful
    • gel 3: 36-6 successful
    • gel 4: 62-1 & 62-3 successful

Thursday 8/12

  • Prepped 4 successful (non-red) cultures: 35-2, 36-6, 62-1, 63-3
  • Digested:
    • 35 (us)
    • 36 (us)
    • 37 (ds)
    • 63 (us)
  • Started cultures of 62-1, BBa_I15008, BBa_I15009
  • Ligations:
    • L76: L35 + L20 + Tet
    • L58: L36 + L37 + Cm
    • L43: L17 + HSP-amp + Tet
  • CPCR: 31, 32, 33, 34, 38, 41, 49, 60, 61, 64
    • gels appeared to have no bands at all... will re-run tomorrow.
  • Transformed L76, L58, L43

Friday 8/13

  • Ligation results:
    • L76: 0 colonies; failed
    • L58: > 500 colonies
    • L43: ~100 colonies
  • CPCR gels rerun from yesterday appear just as bad. 64-5 looks correct, but all other lanes are empty.
  • Cultured L55-5
  • Digested: BBa_I15008 (ds), BBa_I15009 (ds), L42 (bbone)
  • Ligated:
  • CPCR: L76 x 6, L58 x 3; will run the gels tomorrow.
  • Transformed: L31, L32, L33, L34, L41, L60, L61, L76, L77, L78, L79, L80, BBa_I15010
  • AIBN crosslinking ability test with linseed oil and hydrolized linseed oil.
    • 3%-wt of AIBN mixed with 20ml of solution
    • React for 5 hours at 60 degress Celsius water bath.
    • Solution left at room at room temperature overnight
    • white rod-shaped crystals were found in the solution second day
    • the crystals are suspected to be products of AIBN self-reaction
    • no sign of crosslinking

Weekend 8/14-8/15

Saturday 8/14

  • Ligation results:
    • Success: L31, L34, L77, L78, L79, L80
    • BBa_I15010 and L33 both had < 5 colonies
    • Failed: L32, L76, L41, L60, L61
  • Ran gel of yesterday's CPCR results: no significant bands
  • More CPCR
    • 3x of each: L31, L34, L77, L78, L79, L80, L33, L58, L43


 
Caltech logo.png
Caltech footer.png


Locations of visitors to this page